# Extreme mass distributions for $k$-increasing quasi-copulas

### v1

@@ 4. Proofs of Theorems 1 and 2
5. Proofs of Theorems 1 and 2
Finally we can prove Theorem 1.
Proof of Theorem 1.
Using Claim 1 and (3.11), the linear program (3.19) simplifies to:
$y_{2}^{\ast}=y_{1}^{\ast}+dy_{4}^{\ast}.$
Using Claim 2, (5.1) simplifies to
Clearly, the optimal solution satisfies
which proves the first part of Theorem 1.
Proof of Theorem 2.
The proof is analogous to the proof of Theorem 1. ∎
@@ 6. Concluding remarks and future research
In this paper, we addressed the problem of determining the extreme values of the mass distribution associated with a multidimensional $k$-increasing $d$-quasi-copula for $k=2,3,\ldots,d$, building on the linear programming approach that solved the case $k=1$ (see [6] for $d=3$, [14] for $d=4$, [5] for $d\leq 17$, and [13] for the general case). We derived closed formulas for the extreme volumes as well as an explicit realization of one corresponding quasi-copula. Besides the symmetrization trick from [13], the main novelty enabling us to solve the linear programs was the introduction of new sets of variables, which significantly reduced their complexity. We also provided a numerical analysis of the results for dimensions $d\leq 15$.
Finally, we outline several directions for future research.
In the bivariate case, the 2-increasing property coincides with the supermodularity of a function $F$. Very recently, Anzilli and Durante [2] derived bounds on the average $F$-volume of closed rectangles in $0,1]^{2}$ for a supermodular aggregation function $F$. In line with these results, it would be interesting to study the behavior of extreme volumes for supermodular quasi-copulas in higher dimensions, as well as for ultramodular and modular quasi-copulas investigated in [11].

### v2

@@ 5. Concluding remarks and future research
In this paper, we addressed the problem of determining the extreme values of the mass distribution associated with a multidimensional $k$-increasing $d$-variate quasi-copula for $k=2,3,\ldots,d$, building on the linear programming approach that solved the case $k=1$ (see [8] for $d=3$, [9] for $d=4$, [10] for $d\leq 17$, and [11] for the general case). We derived closed formulas for the extreme volumes as well as an explicit realization of one corresponding quasi-copula. Besides the symmetrization trick from [11], the main novelty enabling us to solve the linear programs was the introduction of new sets of variables, which significantly reduced their complexity. We also provided a numerical analysis of the results for dimensions $d\leq 15$.
Finally, we outline several directions for future research. In the bivariate case, the 2-increasing property coincides with the supermodularity of a function $F$. Very recently, Anzilli and Durante [3] derived bounds on the average $F$-volume of closed rectangles in $0,1]^{2}$ for a supermodular aggregation function $F$. In line with these results, it would be interesting to study the behavior of extreme volumes for supermodular quasi-copulas or supermodular aggregation functions [14] in higher dimensions, as well as for ultramodular and modular quasi-copulas investigated in [12].